Consider an experiment with six groups, with nine values in each. For the ANOVA summary table shown to the right, fill in allWill rate, thank you in advance.
Answer #1

sauce df Ss MSS Ambro 2 o 6-1 = s 20*5=loo 22 480 gx6-6=48 uslupalo within Totel 53 45९० CS Scanned with CamScanner